OpenStudy (anonymous):

Which algebraic expression is equivalent to the expression below? 5(4x + 4) a 40x b 20x + 4 c 20x + 20 d 4x + 20

OpenStudy (anonymous):

@clamin

OpenStudy (anonymous):

@thebro55

OpenStudy (anonymous):

@kelliandco

OpenStudy (anonymous):

@moneybaby980

Nnesha (nnesha):

distribute multiply everything which is in parentheses by outside term for example \[\huge\rm \color{red}{a}(b+c) = \color{red}{a} \times b+ \color{red}{a} \times c\] \[\large\rm =\color{red}{a}b+\color{Red}{a}c\]

Nnesha (nnesha):

inside*